OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify this rational expression - (4x^2-6x-18) (x^2+1) ------------ * ----------- (x+1) (x^2-6x+9)

ganeshie8 (ganeshie8):

Have you tried factoring the quadratics ?

OpenStudy (abbycross167):

Umm I've tried to factor it, but I got stuck at a step, would you like to see the steps I did?

ganeshie8 (ganeshie8):

Sure..

OpenStudy (abbycross167):

@ganeshie8 I was told to mutiply first - (4x^2−6x−18)(x^2+1) --------------------- (x-1)(x^2-6x+9)

ganeshie8 (ganeshie8):

Looks good Next try factoring the quadratics

OpenStudy (abbycross167):

See that is where she left me... We were doing this thing called "box method"

ganeshie8 (ganeshie8):

(4x^2-6x-18) = 2(2x^2 - 3x - 9)

ganeshie8 (ganeshie8):

Pick two numbers that addup to -3 and multiply to -18

OpenStudy (abbycross167):

-14+11=-3

ganeshie8 (ganeshie8):

Do they multiply to -18 too ?

OpenStudy (abbycross167):

No sir/ma'am

ganeshie8 (ganeshie8):

So -14 and 11 don't work Pick some other numbers

OpenStudy (abbycross167):

Give me another minute, I'm not very good with negative numbers

ganeshie8 (ganeshie8):

Take your time This is your chance to master them

OpenStudy (abbycross167):

Ok so -6 and 3 -6+3=-3 -6*3=18

ganeshie8 (ganeshie8):

Perfect! so -6 and 3 are your magic numbers

ganeshie8 (ganeshie8):

(4x^2-6x-18) = 2(2x^2 `- 3x` - 9) split the middle term `-3x` as `-6x + 3x`

ganeshie8 (ganeshie8):

(4x^2-6x-18) = 2(2x^2 `- 3x` - 9) = 2(2x^2 `- 6x + 3x` - 9)

ganeshie8 (ganeshie8):

Convince yourself that they both are same

ganeshie8 (ganeshie8):

Next, group first two terms and last two terms : (4x^2-6x-18) = 2[2x^2 `- 3x` - 9] = 2[2x^2 `- 6x + 3x` - 9] = 2[(2x^2 `- 6x`) + ( `3x` - 9)]

ganeshie8 (ganeshie8):

Now factor out the GCF from each group : (4x^2-6x-18) = 2[2x^2 `- 3x` - 9] = 2[2x^2 `- 6x + 3x` - 9] = 2[(2x^2 `- 6x`) + ( `3x` - 9)] = 2[2x(x - 3) + 3( x - 3)]

ganeshie8 (ganeshie8):

Lastly factor the common x-3 : (4x^2-6x-18) = 2[2x^2 `- 3x` - 9] = 2[2x^2 `- 6x + 3x` - 9] = 2[(2x^2 `- 6x`) + ( `3x` - 9)] = 2[2x(x - 3) + 3( x - 3)] = 2[(x - 3)(2x + 3)]

ganeshie8 (ganeshie8):

Still wid me ?

OpenStudy (abbycross167):

Yes sir/ma'am

ganeshie8 (ganeshie8):

(4x^2−6x−18)(x^2+1) --------------------- (x-1)(x^2-6x+9)

ganeshie8 (ganeshie8):

plugin the factored form for the top quadratic

ganeshie8 (ganeshie8):

(4x^2−6x−18)(x^2+1) --------------------- (x-1)(x^2-6x+9) is same as 2(x - 3)(2x + 3)(x^2+1) --------------------- (x-1)(x^2-6x+9)

ganeshie8 (ganeshie8):

see if you can factor the bottom (x^2-6x+9) similarly

ganeshie8 (ganeshie8):

pick two numbers that add up to -6 and multiply to +9

OpenStudy (abbycross167):

@ganeshie8 I can't figure this one out

ganeshie8 (ganeshie8):

How about -3 and -3 ?

ganeshie8 (ganeshie8):

x^2 - 6x + 9 = x^2 - 3x - 3x + 9 = ?

OpenStudy (abbycross167):

Oh they can be the same number?!

OpenStudy (abbycross167):

x^2-3-3+9=4

OpenStudy (abbycross167):

@ganeshie8

ganeshie8 (ganeshie8):

No. You cannot add apples and oranges

ganeshie8 (ganeshie8):

Don't neglect x

ganeshie8 (ganeshie8):

The idea here is to group and factor

ganeshie8 (ganeshie8):

x^2 - 6x + 9 = x^2 - 3x - 3x + 9 group first two terms and last two terms : = (x^2 - 3x) + (- 3x + 9)

ganeshie8 (ganeshie8):

whats the GCF in first group ?

OpenStudy (abbycross167):

1? I'm not sure